OpenStudy (anonymous):

David wants to hang a mirror in his room but the mirror and frame must not have an area larger than 8 square feet. The mirror is 2 feet wide and 3 feet long. Which quadratic equation represents the area of the mirror and frame combined? (The frame must have equal width of x on each side.)

OpenStudy (anonymous):

A:2x^2+14x-2=0 B:3x^2+10x-8=0 C:4x^2+10x-2=0 D:x^2+7x-8=0

OpenStudy (anonymous):

The answer is C because \[(2x+2)(2x+3)=8\rightarrow4^2+10x+6=8\] \[4x^2+10x-2=0\]